# Privacy on the Fly: A Predictive Adversarial Transformation Network for Mobile Sensor Data

### v2

@@ Introduction
Mobile applications can easily access motion sensor data (e.g., accelerometer and gyroscope) through APIs on Android and iOS platforms (rajguru2019sensor). This supports various functions like activity recognition (wang2019deep), step counting (susi2013motion), and gesture interaction (brotchie2022leveraging), making motion data central to mobile services. However, this widespread access also raises privacy concerns, as sensor data can reveal personal attributes like identity (fereidooni2023authentisense), gender (meena2020gender), and age (lin2023childshield). Moreover, the availability of open-source privacy inference models (malekzadeh2019motionsense) makes it easier for third-party applications to exploit such data, potentially infringing on user privacy without their knowledge. These concerns underscore the need for mechanisms that can safeguard user privacy without compromising the utility of motion sensor data.
@@
Adversarial perturbations exploit the sensitivity of machine learning models by adding imperceptible noise to inputs, causing incorrect outputs. We adopt such perturbations as a lightweight strategy to suppress sensitive patterns while preserving the temporal semantics of sensor data. While adversarial attacks on time-series classification models (yang2022tsadv; pialla2025time) have been widely studied, most operate on complete input segments, making them unsuitable for streaming scenarios and unable to address limitation (2). Simply applying perturbations generated from historical data directly to future inputs has limited effectiveness, as it may fail to align with the temporal dynamics of the target sequence. Universal adversarial perturbations (rathore2020untargeted) often perform poorly on complex, high-dimensional sensor data, further limiting their practicality.
@@
To protect user privacy while preserving sensor data semantics, we propose the PATN framework, as shown in Figure 3. Assuming access to open-source privacy inference models and their gradients (white-box), PATN leverages historical sensor data to predict future-directed perturbations, optimized to mislead privacy inference models with minimal semantic distortion. The framework has two components: (1) PATN training, where PATN is optimized using three objectives—adversarial effectiveness, temporal robustness, and smoothness regularization—to balance privacy protection and data fidelity; and (2) On-device deployment, where the trained network is executed within a Trusted Execution Environment (TEE) to enable secure, Zero-latency perturbation generation for real-time sensor streams.
@@
In this section, we present the experimental settings and evaluate our method through baseline comparisons. We conduct utility analysis to examine the trade-off between privacy protection and task performance, perform ablation studies to assess each component’s contribution, and carry out black-box migration studies to test robustness.
@@
Most privacy inference models for sensor data adopt CNN architectures. In this work, we examine two representative CNN-based models for extracting sensitive user information, following the architectures of Lin et al. (lin2023childshield) (age) and Sharshar et al. (sharshar2021activity) (gender). We also evaluate the performance of PATN on sequential models (RNN, LSTM) and report the results in the Appendix (PATN2025).
@@
To further evaluate the transferability of PATN, we evaluate its robustness against three unseen black-box privacy inference models: MobileNet, Xception, and a fully convolutional network (FCN). Although PATN is trained in a white-box setting with three CNNs (see Section Multiple Privacy Inference Models.: Multiple Privacy Inference Models), these models introduce entirely different architectures. As shown in Table 8, PATN maintains consistent performance under these transfer scenarios, achieving ASR values of 29.43%, 36.57%, and 33,48%, and EER of 36.76%, 37.89%, and 38.79% on MobileNet, Xception, and FCN, respectively. These results show that PATN preserves a satisfactory level of privacy protection even against black-box models with unseen or more complex architectures.
@@ Conclusion
In this paper, we propose a history-driven perturbation generation framework PATN, that leverages past sensor data to synthesize future adversarial traces. By integrating a generative network conditioned on historical signals, PATN enables real-time, zero-latency obfuscation while preserving high fidelity with the original data. This design ensures that perturbations can be applied instantly as new data arrive, maintaining strong alignment with natural sensor patterns and effectively protecting sensitive user information from inference-based privacy threats.
For future work, we plan to further investigate the effectiveness of our method in black-box settings, expanding our evaluation to a wider range of model architectures beyond those already studied. Additionally, we aim to explore the generalizability of PATN to a broader spectrum of privacy-sensitive attributes, including but not limited to gender and age.
@@ Acknowledgments
The authors sincerely thank all participants of the peer review process. Special appreciation is extended to Xuanqi Gao, Yuhan Zhi, Weipeng Jiang, Yulong Yang, Yuchen Ren, and Chen Ma for their valuable support and insightful suggestions that contributed to the completion of this work. This work is supported by the National Key Research and Development Program of China (2023YFB3107401), the National Natural Science Foundation of China (T2341003, 62521002, 62376210, 62161160337, 62132011, U24B20185, U21B2018, 62206217), the Shaanxi Province Key Industry Innovation Program (2023-ZDLGY-38), JSPS KAKENHI JP23K24851, JST PRESTO JPMJPR23P5, JST CREST JPMJCR21M2, JST NEXUS JPMJNX25C4. Thanks to the New Cornerstone Science Foundation and the Xplorer Prize.
Codehttps://github.com/skysky4/PATN
Datasetshttps://github.com/mmalekzadeh/motion-sense
APPENDIX
More Details about PATN
In order to protect user privacy on mobile devices, we generate adversarial perturbations for future accelerometer and gyroscope signals by leveraging historical sensor data. The validity of this approach can be justified from three theoretical perspectives.
1) Temporal Correlation and Local Stationarity.
Inertial sensor readings can be modeled as
$x_{t+1}=g(x_{t})+\epsilon_{t}$ (11)
where $g(\cdot)$ denotes a smooth dynamical process and $\epsilon_{t}$ is a noise term. Over short horizons, $\epsilon_{t}$ is approximately weakly stationary, implying
$p(\epsilon_{t+k})\approx p(\epsilon_{t}),\quad k\ll T.$ (12)
Thus, the noise statistics extracted from historical data can approximate the distribution of future perturbations.
2) Direct Perturbation Generation without Explicit Prediction.
Modern generative models (e.g., VAEs, conditional GANs, RNN-based predictors) are capable of learning the conditional distribution $p(x_{t+1}|x_{1:t})$ to synthesize future sensor data. Since our goal is not to reconstruct exact future trajectories but to utilize the underlying statistical properties, we can bypass explicit prediction and directly sample adversarial perturbations from the estimated historical noise distribution:
$\delta_{t}\sim\hat{p}(\epsilon).$ (13)
This avoids compounding prediction errors and yields perturbations more representative of the true signal dynamics.
3) Dynamic Adjustment via Motion Pattern Resampling.
Historical sensor data can be decomposed into characteristic motion patterns $\phi_{i}(t)$. To adapt perturbations to current states, we dynamically combine these patterns:
$\delta_{t}=\sum_{i}w_{i}(t)\phi_{i}(t),$ (14)
where weights $w_{i}(t)$ are updated based on current inputs and optimization goals (e.g., maximizing inference error while controlling perturbation magnitude). This ensures the generated perturbations remain physically plausible while effectively degrading privacy-invasive inference.
Overall, by exploiting the local stationarity of sensor noise, leveraging the statistical modeling capability of generative frameworks, and applying dynamic resampling, we can efficiently construct adversarial perturbations for future sensor readings without explicit signal prediction.
More Details about the Datasets
MotionSense (Malekzadeh:2018:PSD:3195258.3195260) is an open-source dataset that contains data captured from an accelerometer and a gyroscope at a constant frequency of 50 Hz, collected using an iPhone 6s placed in the participants’ front pocket. In total, 24 participants performed six distinct activities—going downstairs, going upstairs, walking, jogging, sitting, and standing—over 15 trials, all conducted in a controlled environment. This dataset is suitable for studying privacy inference of gender based on motion patterns. However, as all participants are adults, it cannot be used for privacy inference related to distinguishing between children and adults.
ChildShield (lin2023childshield) is a proprietary dataset, for which we obtained permission from the authors for research use. It contains motion sensor data collected from accelerometers and gyroscopes at a constant frequency of 60 Hz during gameplay across multiple mobile games. A total of 1,875 participants completed trials on five different games. This dataset enables studies on privacy inference of age attributes (child and adult users). However, as no gender labels were collected, it cannot be used to train or evaluate gender-based privacy inference models.
Interpretable Feature Perturbation via T-SNE Visualization
Figure 6: T-SNE visualization of feature representations under different privacy perturbation methods: (a) Raw Data, (b) DP, (c) PGD, and (d) PATN.
To better interpret the effect of different privacy defense mechanisms on the internal representations of the inference model, we visualize the latent feature space using T-SNE. Specifically, we compare the feature distributions resulting from the raw data, our proposed PATN method, and baseline methods: DP and PGD. These visualizations are based on the hidden representation extracted from the inference model and reflect how input perturbations alter the separability of private attributes in the feature space.
As shown in Figure 6, the raw data exhibits clear clustering patterns aligned with the private attribute labels, indicating that the inference model can easily exploit these structures. In contrast, the PATN method significantly disrupts these patterns, making the latent representations more entangled and less linearly separable. Compared to DP and PGD, PATN achieves a more substantial deformation of the feature space while maintaining naturalness in the perturbed signals. This illustrates the interpretability advantage of PATN: rather than merely injecting noise, it learns to adaptively shift samples in the feature space to confuse the inference model effectively. The visual evidence thus supports the claim that PATN provides stronger and more targeted privacy protection through latent feature perturbation.
Figure 7: Comparison of Raw and Adversarial sensor data (gyroscope) on IMU Axes.
IMU Visualization of Gyroscope Data
To better interpret our method’s effect in the sensor space, we also visualize gyroscope IMU signals, as shown in Figure 7. The gyroscope visualization shows that our perturbations preserve strong semantic fidelity: key motion patterns and temporal dynamics remain intact despite the added noise. This indicates that the method maintains signal realism essential for utility tasks, while effectively masking features used for privacy inference.
Method RNN LSTM
$uparrow ASR()\uparrow EER()\uparrow$
Raw data \ 14.76 \ 13.54
DP 5.05 19.52 1.50 14.56
UAP 2.45 15.92 1.04 14.24
FGSM 8.52 20.90 6.74 18.44
PGD 8.52 20.90 6.11 17.73
PATN(our) 31.59 29.26 25.35 25.01
Table 9: Performance comparison with baseline methods.
Effectiveness of PATN on Temporal Models
To further evaluate the generalizability of PATN, we extend our experiments to sequential models, including RNN and LSTM, using the MotionSense dataset. As shown in Table 9, PATN significantly outperforms all baseline methods in both ASR (Attack Success Rate) and EER (Equal Error Rate), across both model types. Notably, PATN achieves 31.59% ASR and 29.26% EER against the RNN-based inference model, while the best-performing baseline (FGSM/PGD) reaches only 8.52% ASR and 20.90% EER. Similarly, on the LSTM model, PATN attains 25.35% ASR and 25.01% EER, considerably higher than all baselines (e.g., PGD: 6.11% ASR, 17.73% EER).
We observe that attacking sequential models such as RNNs and LSTMs is generally more challenging than attacking CNNs, possibly due to their non-convex optimization landscape and temporally entangled feature representations (ding2023black). While the overall adversarial effectiveness is reduced in this setting, our method still provides a noticeable level of privacy protection, outperforming all baselines in both ASR and EER. These results suggest that PATN maintains a certain degree of robustness across different model architectures, even under complex temporal modeling conditions. In future work, we plan to further investigate specialized perturbation strategies tailored for recurrent models, in order to enhance privacy defense in sequential settings.

### v4

@@ Introduction
Mobile applications can easily access motion sensor data (e.g., accelerometer and gyroscope) through APIs on Android and iOS platforms (rajguru2019sensor). This supports various functions like activity recognition (wang2019deep), step counting (susi2013motion), and gesture interaction (brotchie2022leveraging), making motion data central to mobile services. However, this widespread access also raises privacy concerns, as sensor data can reveal personal attributes like identity (cai2024famos; fereidooni2023authentisense; shen2017performance), gender (meena2020gender), and age (lin2023childshield; miao2023learning). Moreover, the availability of open-source privacy inference models (malekzadeh2019motionsense) makes it easier for third-party applications to exploit such data, potentially infringing on user privacy without their knowledge. These concerns underscore the need for mechanisms that can safeguard user privacy without compromising the utility of motion sensor data.
@@
Adversarial perturbations exploit model sensitivity by adding imperceptible noise that induces incorrect outputs. We adopt them as a lightweight way to suppress sensitive patterns while preserving sensor data semantics. Although adversarial attacks on time-series models (yang2022tsadv; pialla2025time) are well studied, most rely on full input segments, making them unsuitable for streaming and unable to address limitation (2). Simply applying perturbations generated from historical data directly to future inputs has limited effectiveness, as it may fail to align with the temporal dynamics of the target sequence. Universal adversarial perturbations (rathore2020untargeted) often perform poorly on complex, high-dimensional sensor data, further limiting their practicality.
@@
To protect user privacy while preserving sensor data semantics, we propose the PATN framework, as shown in Figure 3. Assuming access to open-source privacy inference models and their gradients (white-box), PATN leverages historical sensor data to predict future-directed perturbations, optimized to mislead privacy inference models with minimal semantic distortion. The framework has two components: (1) PATN training, where PATN is optimized using three objectives—adversarial effectiveness, temporal robustness, and smoothness regularization—to balance privacy protection and data fidelity; and (2) On-device deployment, where the trained network is executed locally on mobile devices to enable zero-latency perturbation generation for real-time sensor streams.
@@
In this section, we describe the experimental setup and evaluate our method via baseline comparisons, utility analysis, ablation studies, and black-box migration tests to assess performance, component contributions, and robustness.
@@
Most privacy inference models for sensor data adopt CNN architectures. In this work, we examine two representative CNN-based models for extracting sensitive user information, following the architectures of Lin et al. (lin2023childshield) (age) and Sharshar et al. (sharshar2021activity) (gender). We also evaluate PATN on sequential models (RNN, LSTM), with results reported in the Appendix (PATN2025).
@@
To further evaluate PATN’s transferability, we test its robustness against three unseen black-box privacy inference models: MobileNet, Xception, and a fully convolutional network (FCN). Although trained in a white-box setting with three CNNs (Section Multiple Privacy Inference Models.), these models have entirely different architectures. As shown in Table 8, PATN maintains consistent performance, achieving ASR of 29.43%, 36.57%, and 33.48%, and EER of 36.76%, 37.89%, and 38.79% on MobileNet, Xception, and FCN, respectively, demonstrating satisfactory privacy protection even against unseen or more complex models.
@@ Conclusion
In this paper, we propose PATN, a history-driven perturbation generation framework that leverages past sensor data to synthesize future adversarial traces. By conditioning a generative network on historical signals, PATN enables real-time, zero-latency obfuscation while preserving fidelity to the original data. Perturbations can be applied instantly to incoming data, maintaining alignment with natural sensor patterns and mitigating inference-based privacy risks. For future work, we plan to evaluate PATN in black-box settings across more model architectures and extend its applicability to additional privacy-sensitive attributes, including but not limited to gender and age.
@@ Acknowledgments
The authors sincerely thank all participants of the peer review. Special appreciation is extended to Xuanqi Gao, Yuhan Zhi, Weipeng Jiang, Yulong Yang, Yuchen Ren, and Chen Ma for their valuable support and insightful suggestions that contributed to the completion of this work. This work is supported by the National Key Research and Development Program of China (2023YFB3107401), the National Natural Science Foundation of China (62402377, 62536002, T2341003, 62521002, 62376210, 62161160337, 62132011, U24B20185, U21B2018, 62206217), the Shaanxi Province Key Industry Innovation Program (2023-ZDLGY-38), JSPS KAKENHI JP23K24851, JST PRESTO JPMJPR23P5, JST CREST JPMJCR21M2, JST NEXUS JPMJNX25C4. Thanks to the New Cornerstone Science Foundation and the Xplorer Prize.